Reese, a calendar-year taxpayer, uses the cash method of accounting for her sole proprietorship. In late...

  1. Reese, a calendar-year taxpayer, uses the cash method of accounting for her sole proprietorship. In late December she received a $30,000 bill from her accountant for consulting services related to her small business. Reese can pay the $30,000 bill any time before January 30 of next year without penalty. Assume Reese’s marginal tax rate is 24 percent this year and will be 32 percent next year, and that she can earn an after-tax rate of return of 8 percent on her investments. When should she pay the $25,000 billthis year or next?

    December

    January

    You should be indifferent because the net cost is the same in either case.

    You would need additional information to determine whether the payment should be made in December or January.

QUESTION 6

  1. Reese, a calendar-year taxpayer, uses the cash method of accounting for her sole proprietorship. In late December she received a $30,000 bill from her accountant for consulting services related to her small business. Reese can pay the $30,000 bill any time before January 30 of next year without penalty. Assume Reese’s marginal tax rate is 24 percent this year and will be 22 percent next year, and that she can earn an after-tax rate of return of 8 percent on her investments. When should she pay the $30,000 billthis year or next?

    December

    January

    Indifferent as the net benefit is the same in both cases.

    You need additional information to make a recommendation.

QUESTION 7

  1. Assume Rafael can earn an 6 percent after-tax rate of return. He can receive $2,000 today or $3,000 in five years. What is the present value of the $3,000 payment due in five years?

QUESTION 8

  1. Assume Ellina earns a 8 percent after-tax rate of return, and that she owes a friend $2,400. She can pay the friend $2,400 today or $3,500 in four years. What is the present value of the $3,500 payment due in four years?
Answer #1

SOLUTION OF QUES 6-

AMOUNT PAY IN DECEMBER

30000*24%= $ 7200

30000-7200= $ 22800

AMOUNT PAY IN JANUARY

30000*22%= $ 6600

PRESENT VALUE OF TAX=6600 *0.926(DISCOUNTING FACTOR) (1 YEAR ,8%)

= $ 6111.6

AFTER TAX INCOME =Pretax income-present value tax

=30000-6111.6

= $ 23888.4

SOLUTION OF QUES 7-

PRESENT VALUE OF THE $ 3000 IS- 3000*0.747 (DISCOUNTING FACTOR OF 5 YEAR ,RATE 6%

= $ 2241

SOLUTION OF QUES 8-

TODAY = $ 2400

$ 3500 IN FOUR YEARS IS WORTH TODAY = 3500 * 0.735 (DISCOUNT FACTOR OF FOUR YEAR,8%)

= $ 2572.5

SO PAYING TODAY $ 2400 IS MORE PREFERABLE THAN PAYING $ 3500 (WORTH $ 2572.5 TODAY) IN FOUR YEARS.